OpenStudy (anonymous):

A cat falls from a table of height 1.3 m. What is the impact speed of the cat?

OpenStudy (anonymous):

assume that the initial velocity is zero use the formula vf^2=vi^2+2ad where a=9.8m/s^2 and vi=0m/s and d=1.3m